# CO-VADA: A Confidence-Oriented Voice Augmentation Debiasing Approach for Fair Speech Emotion Recognition

### v1

@@ Abstract
Index Terms:
Speech emotion recognition, fairness, data augmentation, voice conversion, responsible and trustworthy machine
@@ I Introduction
However, to the best of our knowledge, very few (if any) studies have explored debiasing SER systems without relying on demographic information. Existing efforts in related speech tasks primarily focus on model-based interventions, such as adversarial training or sample reweighting [10, 11, 12], which often require architectural modifications and may lack generalizability. In contrast, data-centric approaches (i.e., data augmentation or distribution balancing) tend to offer broader compatibility across models [13, 14]. Nevertheless, most of these methods still depend on explicit demographic annotations (e.g., gender, age, or speech disorder) to identify and mitigate bias. While these techniques have not been widely applied to SER, their general strategies hold promise for bias mitigation in this domain. Unfortunately, in real-world applications, demographic metadata is often unavailable, incomplete, or unreliable. These limitations underscore the need for effective debiasing techniques that operate independently of demographic information.
@@ II-B Debiasing Methods with Demographic Information
Recent debiasing methods for speech-input tasks rely on a variety of techniques, including counterfactual data augmentation through voice conversion [26], cross-lingual adaptation combined with SpecAugment and speed perturbation [14, 27], and subgroup-aware sampling guided by divergence metrics [28, 29]. Besides, other approaches include group-adapted fusion networks for fairer speaker verification (SV) [30], adversarial training [31, 32], gender neutralization [18], and contrastive learning to enhance subgroup-invariant representations in intent classification [33]. While these techniques have successfully reduced demographic performance gaps in ASR, SV, and Spoken Language Understanding tasks, they rely on predefined demographic labels, which limit their applicability to unannotated or intersectional biases.
@@ II-C Debiasing Methods without Demographic Information
Additionally, SiH [36] utilizes a Variational Autoencoder (VAE) combined with focal-loss reweighting to develop both biased and unbiased classifiers, jointly down-weighting spurious patterns. BLIND [37] trains a “success detector” that predicts when the main model’s decisions hinge on simplistic, shortcut features, subsequently down-weighting these samples through a debiased focal loss [38], thereby dynamically alleviating bias using only task labels.
@@ III-A Sample Separation by Confidence (Step 1-3)
To analyze this potential bias further, we define the emotion-specific subsets of the data. Let E denote the set of all emotion labels in the dataset. For each emotion category $e\in E$, we extract the label subset:
@@ IV-B True Positive Rate Gap
To assess fairness in classification performance across demographic groups, we compute $\mathrm{TPR}_{\text{gap}}$ [49]. Let $\mathcal{Z}$ denote the set of demographic groups (e.g., genders), and $E$ the set of emotion classes. For each class $e\in E$ and group $z\in\mathcal{Z}$, the true positive rate is defined as $\mathrm{TPR}_{e}^{(z)}$.
@@ IV-C Demographic Parity Gap
While it might seem logical to normalize this rate by the specific sample counts for each class, it’s important to note that in a multi-label setting, every sample can be associated with all classes. Therefore, we consistently use the total number of samples $N$ for all classes.
@@ V-A Model Setup
In our experiments, we used WavLM-base+  [^1] https://github.com/microsoft/unilm/tree/master/wavlm [51] as the feature extractor. Instead of relying only on the final layer, we applied a learnable weighted sum over all feature layers to obtain the final representation [52]. This representation was then passed on to a simple two-layer linear classifier. We trained all models using AdamW [53] optimizer with a learning rate of 1e-4 and a batch size of 32.
@@ V-B Datasets
While demographic labels are not used during model training, we utilized available metadata during evaluation to compute fairness metrics. The three datasets differ in their demographic coverage and the scope of fairness analysis. The CREMA-D included annotations for gender, age, and race, enabling evaluation on multiple demographic axes. The MSP-Podcast and MSP-IMPROV only provided gender metadata, thus fairness evaluation on these datasets was restricted to gender-based analysis. In addition, the CREMA-D and MSP-IMPROV were partitioned into five and six folds, respectively, following a cross-validation setting in the EMO-SUPERB [50].
@@ V-C Voice Conversion Models
VC was used in our framework as a critical tool to generate speaker-varied audio samples for data augmentation. We primarily adopted FreeVC [58] for all main experiments, due to its high-quality, text-free, one-shot conversion capability using WavLM features.
To evaluate the generality of our framework across various VC implementations, we included two additional models in our ablation studies: Diff-HierVC [59], which is a diffusion-based model featuring a hierarchical architecture, and kNN-VC [60], a nonparametric method that performs frame-level matching without the need for training.
We directly applied all VC models in their pretrained form, without any fine-tuning or model-specific optimization. This design enabled us to assess the robustness of our debiasing approach and confirmed that its effectiveness did not rely on a particular voice conversion model.
@@ V-D Bias Partitioning and Augmentation Strategy
To identify samples that reflect biases at the dataset level, we first trained an initial emotion classifier with early stopping. The main purpose is to capture the model’s early learning dynamics, a phase in which it tends to rely on spurious correlations. Early stopping was triggered once the Macro F1 exceeds a predefined threshold of 0.5 for the CREMA-D and MSP-IMPROV, and 0.3 for the MSP-Podcast. The lower threshold for MSP-Podcast reflected the particular difficulty in achieving high performance in SER using this dataset. In our experiments, model performance on the MSP-Podcast rarely exceeds a Macro F1 of 0.5 during training.
We adopted a consistent strategy for pairing and sample generation. Within each emotion category, training samples were ranked in ascending order by cross-entropy loss. The lower half of the ranked samples was designated as the bias-guiding set, while the upper half formed the bias-contrary set. We then created random one-to-one pairings between the two sets. The number of pairs per emotion matched the number of original samples in that category.
@@ V-E Baselines and Selection Criteria
We compared CO-VADA with six representative de-biasing baselines: LVR [34], SiH [36], BLIND [37], IDI-RW [39], DisEnt [35], and LfF [15]. These methods cover a range of debiasing strategies, including reweighting, disentangled representation learning, adversarial training, and selective sampling. We focused on approaches that do not rely on demographic labels, ensuring consistency with our training setting. To allow a fair comparison, all methods were evaluated under the same experimental configuration. The ERM was the original baseline without any debiasing method, following [39].
@@ VI Experimental Results
The proposed CO-VADA achieved the best balance between classification performance and fairness across gender, race, and age based on results from the CREMA-D dataset. Fig. 2 illustrates these results in performance-fairness scatter plots. It appeared consistently in the lower right region of the F1–$\mathrm{TPR}_{\text{gap}}$ and F1–$\mathrm{DP}_{\text{gap}}$ scatter plots, indicating strong predictive accuracy with low subgroup disparities. Among the baselines, SiH, DisEnt, and LfF reached similar Macro F1 but exhibited substantially larger fairness gaps, suggesting that these methods did not effectively mitigate bias. Moreover, LVR performed comparably with ours in fairness, particularly for gender, but suffered from a noticeable drop in Macro F1. IDI-RW approached our fairness levels, yet its classification performance remained consistently lower. ERM, which does not employ a debiasing strategy, yielded moderate accuracy but poor fairness.
Regarding the results on the MSP-Podcast, our method achieved the lowest $\mathrm{TPR}_{\text{gap}}$ among all baselines. Fig. 3 shows the corresponding plots for gender fairness. These findings demonstrated that our method achieved the most consistent recognition of emotions between gender subgroups, which is particularly important in fairness-sensitive applications. While our Macro F1 was slightly lower than LVR’s, the difference was negligible. LVR and our method demonstrated comparable $\mathrm{TPR}{\text{gap}}$ and $\mathrm{DP}{\text{gap}}$ values, indicating similar fairness performance across demographic groups. Besides, SiH achieved a marginal improvement in Macro F1, but exhibited substantial fairness gaps. LfF also did not match our method in either metric.
@@
Table II: Effect of macro F1 thresholds used for early stopping. Each row shows the results of running the full debiasing pipeline using a different threshold. $\uparrow$ indicates higher is better (SER performance), while $\downarrow$ indicates lower is better (fairness).
Threshold Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$downarrow $\boldsymbol{\mathrm{DP}_{\text{gap}}}$downarrow
0.3 0.653 0.231 0.088
0.4 0.658 0.245 0.089
0.5 0.653 0.233 0.082
0.6 0.655 0.235 0.090
0.7 0.653 0.233 0.087
@@ VII-A Sensitivity of Model Performance to Early Stopping Thresholds
To evaluate the effect of early stopping on bias mitigation, we varied the Macro F1 threshold used to trigger early termination during training on the CREMA-D. The thresholds ranged from 0.3 to 0.7.
The results in Table II showed minimal changes in classification performance between thresholds, with Macro F1 remaining stable. Fairness metrics exhibited slight variations, but the differences were small. This indicated that the specific choice of the early stopping threshold has a limited influence. It did not substantially affect the model’s ability to detect or reduce bias under the current framework.
@@ VII-B Impact of Bias Proportion in Training Data
We evaluated how different proportions of bias-guiding and bias-contrary samples influence model performance and fairness. As shown in Table III, all configurations yielded comparable Macro F1, suggesting that classification performance was largely unaffected by the sampling ratio. However, slight but consistent differences appeared in $\mathrm{TPR}_{\text{gap}}$ and $\mathrm{DP}_{\text{gap}}$. This pattern suggested that fairness was more sensitive than accuracy to the composition of augmented samples.
Table III: Impact of bias-contrary:unused:bias-guiding ratios on performance and fairness. Each ratio represents a partition of samples sorted by cross-entropy loss, with the highest-loss samples used as bias-contrary, the lowest-loss samples as bias-guiding, and the middle portion excluded from augmentation.
Ratio Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$downarrow $\boldsymbol{\mathrm{DP}_{\text{gap}}}$downarrow
3 : 4 : 3 0.654 0.230 0.087
4 : 2 : 4 0.655 0.247 0.092
5 : 0 : 5 0.653 0.233 0.082
4 : 0 : 6 0.653 0.240 0.088
3 : 0 : 7 0.649 0.227 0.083
@@ VII-C Effectiveness of Different Voice Conversion Models
As shown in Table IV, FreeVC achieved the best overall balance between performance and fairness. It maintained high classification accuracy while keeping subgroup disparities low. Diff-HierVC resulted in the lowest fairness gaps, indicating strong bias mitigation, although with a slight reduction in Macro F1. kNN-VC also improved fairness over the ERM baseline but shows lower accuracy, reflecting the limitations of its nonparametric design.
Table IV: SER Performance and fairness under our framework using different VC models.
Methods VC Model Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$downarrow $\boldsymbol{\mathrm{DP}_{\text{gap}}}$downarrow
ERM 0.651 0.278 0.103
CO-VADA kNN-VC[60] 0.642 0.244 0.097
CO-VADA Diff-HierVC[59] 0.648 0.223 0.081
CO-VADA FreeVC[58] 0.653 0.233 0.082
Table V: SER Performance and fairness under different combinations of bias selection (BS) and voice conversion (VC). ✓ indicates the component is included.
# BS VC Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$downarrow $\boldsymbol{\mathrm{DP}_{\text{gap}}}$downarrow
1 0.651 0.278 0.103
2 0.659 0.259 0.097
3 0.652 0.240 0.088
4 0.653 0.233 0.082
@@ VII-D Effect of Bias Selection and Voice Conversion
We compared four configurations that vary in the use of bias selection (BS) and voice conversion (VC) augmentation. Table V summarizes the results.
  • Configuration #1 served as the baseline, which applied neither BS nor VC.
  • Configuration #2 included BS but not VC. In this setting, bias-contrary samples were assigned higher weights during training, but no data were augmented.
  • Configuration #3 included VC but not BS. VC was applied to randomly selected pairs of utterances without considering their bias status.
  • Configuration #4, which represents our complete method, applied both BS and VC by selecting bias-guiding and bias-contrary samples and generating converted examples accordingly.
As shown in Table V, our method (#4) achieved the best overall balance between classification performance and fairness. It produced the lowest $\mathrm{TPR}_{\text{gap}}$ and $\mathrm{DP}_{\text{gap}}$ scores, with competitive Macro F1. Configuration #2 improved the accuracy slightly but offered a limited fairness improvement. Configuration #3 performed better than the baseline in fairness, but its overall results remained inferior to ours.

### v2

@@ I Introduction
However, to the best of our knowledge, very few studies have explored debiasing SER systems without relying on demographic information. Existing efforts in related speech tasks primarily focus on model-based interventions, such as adversarial training or sample reweighting [10, 11, 12], which often require architectural modifications and may lack generalizability. In contrast, data-centric approaches (i.e., data augmentation or distribution balancing) tend to offer broader compatibility across models [13, 14]. Nevertheless, most of these methods still depend on explicit demographic annotations (e.g., gender, age, or speech disorder) to identify and mitigate bias. While these techniques have not been widely applied to SER, their general strategies hold promise for bias mitigation in this domain. Unfortunately, in real-world applications, demographic metadata is often unavailable, incomplete, or unreliable. These limitations underscore the need for effective debiasing techniques that operate independently of demographic information.
@@
Unlike prior debiasing methods in computer vision that rely on class activation maps or spatial structure to localize bias-inducing regions [16], CO-VADA tackles fundamentally different challenges in speech, where such spatial cues are unavailable. Instead of modifying visual content, we operate in the audio domain by altering speaker identity to diversify training data. This targeted augmentation enables debiasing without demographic labels or architectural changes.
@@ II-B Debiasing Methods with Demographic Information
Recent debiasing methods for speech-input tasks rely on a variety of techniques, including counterfactual data augmentation through voice conversion [26], cross-lingual adaptation combined with SpecAugment and speed perturbation [14, 27], and subgroup-aware sampling guided by divergence metrics [28, 29]. Besides, other approaches include group-adapted fusion networks for fairer speaker verification (SV) [30], adversarial training [kim2025domain, 32], gender neutralization [18], and contrastive learning to enhance subgroup-invariant representations in intent classification [33]. While these techniques have successfully reduced demographic performance gaps in ASR, SV, and Spoken Language Understanding tasks, they rely on predefined demographic labels, which limit their applicability to unannotated or intersectional biases.
@@ II-C Debiasing Methods without Demographic Information
Additionally, SiH [vandenhirtz2023signal] utilizes a Variational Autoencoder (VAE) combined with focal-loss reweighting to develop both biased and unbiased classifiers, jointly down-weighting spurious patterns. BLIND [37] trains a “success detector” that predicts when the main model’s decisions hinge on simplistic, shortcut features, subsequently down-weighting these samples through a debiased focal loss [38], thereby dynamically alleviating bias using only task labels.
@@ III-A Sample Separation by Confidence (Step 1-3)
To analyze this potential bias, we define the emotion-specific subsets of the data. Let E be the set of all emotion labels in the dataset. For each emotion category $e\in E$, we extract the label subset:
@@ IV-B True Positive Rate Gap
To assess fairness in classification performance across demographic groups, we compute $\mathrm{TPR}_{\text{gap}}$ [49]. Let $\mathcal{Z}$ denote the set of demographic groups, and $E$ the set of emotion classes. For each class $e\in E$ and group $z\in\mathcal{Z}$, the true positive rate is defined as $\mathrm{TPR}_{e}^{(z)}$.
@@ IV-C Demographic Parity Gap
Although normalizing by class-specific sample counts may seem reasonable, it’s important to note that in a multi-label setting, every sample can be associated with all classes. Therefore, we consistently use the total number of samples $N$ for all classes.
@@ V-A Model Setup
In our experiments, we used WavLM-base+  [^1] https://github.com/microsoft/unilm/tree/master/wavlm [51] as the feature extractor. We applied a learnable weighted sum over all feature layers to obtain the final representation [52], then passed it on to a simple two-layer linear classifier. We trained all models using AdamW [53] optimizer with a learning rate of 1e-4 and a batch size of 32.
@@ V-B Datasets
While demographic labels are not used during model training, we utilized available metadata during evaluation to compute fairness metrics. The three datasets differ in their demographic coverage and the scope of fairness analysis. The CREMA-D included annotations for gender, age, and race, enabling evaluation on multiple demographic axes. The MSP-Podcast and MSP-IMPROV only provided gender metadata; thus, fairness evaluation on these datasets was restricted to gender-based analysis. In addition, CREMA-D was partitioned into five folds and MSP-IMPROV into six folds, respectively, following the cross-validation setting in EMO-SUPERB [50].
@@ V-C Voice Conversion Models
VC was used in our framework to generate speaker-varied audio samples for data augmentation. We primarily adopted FreeVC [58] for all main experiments, due to its high-quality, text-free, one-shot conversion capability using WavLM features.
To evaluate the generality of our framework across various VC implementations, we included two additional models in our ablation studies: Diff-HierVC [59], which is a diffusion-based model featuring a hierarchical architecture, and kNN-VC [60], a nonparametric method that performs frame-level matching without training. We directly applied all VC models in their pretrained form, without any fine-tuning or model-specific optimization.
@@ V-D Bias Partitioning and Augmentation Strategy
To identify samples that reflect biases at the dataset level, we first trained an initial emotion classifier with early stopping. The main purpose is to capture the model’s early learning dynamics, a phase in which it tends to rely on spurious correlations. Early stopping was triggered once the Macro F1 exceeded a predefined threshold of 0.5 for the CREMA-D and MSP-IMPROV, and 0.3 for the MSP-Podcast. The lower threshold for MSP-Podcast reflected the particular difficulty in achieving high performance in SER using this dataset. In our experiments, model performance on the MSP-Podcast rarely exceeds a Macro F1 of 0.5 during training.
We adopted a consistent strategy for pairing and sample generation. Within each emotion category, training samples were ranked in ascending order by cross-entropy loss. The lower half of the ranked samples was designated as the bias-guiding set, while the upper half formed the bias-contrary set. We then created random one-to-one pairings between the two sets. For each emotion category, we generated a number of augmented samples equal to the count of original samples in that category.
@@ V-E Baselines and Selection Criteria
We compared CO-VADA with six representative de-biasing baselines: LVR [34], SiH [vandenhirtz2023signal], BLIND [37], IDI-RW [39], DisEnt [35], and LfF [15]. These methods cover a range of debiasing strategies, including reweighting, disentangled representation learning, adversarial training, and selective sampling. We focused on approaches that do not rely on demographic labels, ensuring consistency with our training setting. To allow a fair comparison, all methods were evaluated under the same experimental configuration. The ERM was the original baseline without any debiasing method, following [39].
To ensure a fair and comprehensive comparison, we evaluated each method under five hyperparameter settings, avoiding selective tuning that could bias the comparison. For each method, we selected one key tunable parameter and tested five different values.
@@ VI Experimental Results
[…]
Regarding the results on the MSP-Podcast, our method achieved the lowest $\mathrm{TPR}_{\text{gap}}$ among all baselines. Fig. 3 shows the corresponding plots for gender fairness. These findings demonstrated that our method achieved the most consistent recognition of emotions between gender subgroups, which is particularly important in fairness-sensitive applications. While our Macro F1 was slightly lower than LVR’s, the difference was negligible. LVR and our method demonstrated comparable $\mathrm{TPR}_{\text{gap}}$ and $\mathrm{DP}_{\text{gap}}$ values, indicating similar fairness performance across demographic groups. Besides, SiH achieved a marginal improvement in Macro F1, but exhibited substantial fairness gaps. LfF also did not match our method in either metric.
@@
Table II: Performance under different macro F1 thresholds used for early stopping. Each row shows the full debiasing pipeline result using a different threshold. $\uparrow$ indicates higher is better, while $\downarrow$ indicates lower is better.
Threshold Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$ $\downarrow$ $\boldsymbol{\mathrm{DP}_{\text{gap}}}$ $\downarrow$
0.3 0.653 +0.31% 0.231 -16.61% 0.088 -14.56%
0.4 0.658 +1.07% 0.245 -11.55% 0.089 -14.56%
0.5 0.653 +0.31% 0.233 -15.88% 0.082 -20.39%
0.6 0.655 +0.61% 0.235 -15.16% 0.090 -12.62%
0.7 0.653 +0.31% 0.233 -15.88% 0.087 -15.53%
@@ VII-A Sensitivity of Model Performance to Early Stopping Thresholds
To evaluate the effect of early stopping on bias mitigation, we varied the Macro F1 threshold used to trigger early termination during training on the CREMA-D. The thresholds ranged from 0.3 to 0.7. The results in TableII show the performance of CO-VADA under each threshold, with subscripts indicating the relative improvements over the baseline ERM (green boxes denote improvements over ERM; red boxes denote degradations). Macro F1 consistently improved across all thresholds, and fairness metrics also showed reductions relative to ERM, though the extent of improvement varied slightly across thresholds. These findings indicate that the specific choice of the early stopping threshold has a limited effect on overall performance and fairness. CO-VADA remains effective in mitigating bias even when the threshold is coarsely selected.
@@ VII-B Impact of Bias Proportion in Training Data
We evaluated how different proportions of bias-guiding and bias-contrary samples influence model performance and fairness. As shown in Table III, all configurations yielded comparable Macro F1, suggesting that classification performance was unaffected by the sampling ratio. However, slight but consistent differences appeared in $\mathrm{TPR}_{\text{gap}}$ and $\mathrm{DP}_{\text{gap}}$. This pattern suggested that fairness was more sensitive than accuracy to the composition of augmented samples.
Table III: Performance under different bias-contrary:unused:bias-guiding ratios. Subscripts indicate relative improvements over the ERM baseline.
Ratio Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$ $\downarrow$ $\boldsymbol{\mathrm{DP}_{\text{gap}}}$ $\downarrow$
3 : 4 : 3 0.654 +0.46% 0.230 -16.96% 0.087 -15.53%
4 : 2 : 4 0.655 +0.61% 0.247 -10.83% 0.092 -10.68%
5 : 0 : 5 0.653 +0.31% 0.233 -15.88% 0.082 -20.39%
4 : 0 : 6 0.653 +0.31% 0.240 -13.36% 0.088 -14.56%
3 : 0 : 7 0.649 -0.31% 0.227 -18.05% 0.083 -19.42%
@@ VII-C Effectiveness of Different Voice Conversion Models
We evaluated the effectiveness of CO-VADA using different VC models. As shown in Table IV, FreeVC achieved the best overall balance in relative improvement over the ERM baseline. It yielded the highest F1 gain while substantially reducing both TPRgap and DPgap. Diff-HierVC achieved the largest reduction in fairness gaps, though with a slight drop in F1. kNN-VC also improved fairness compared to ERM, but showed a relative decrease in classification performance, possibly due to its nonparametric nature.
While the degree of improvement varied depending on the generation quality of each VC model, all models consistently enhanced fairness. This confirms that the effectiveness of our debiasing approach stems from the overall design of CO-VADA, rather than reliance on any particular VC model.
Table IV: SER performance and fairness using different VC models. Subscripts indicate relative improvement over the ERM baseline.
VC Model Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$ $\downarrow$ $\boldsymbol{\mathrm{DP}_{\text{gap}}}$ $\downarrow$
kNN-VC[60] 0.642 -1.38% 0.244 -12.23% 0.097 -5.83%
Diff-HierVC[59] 0.648 -0.46% 0.223 -19.78% 0.081 -21.36%
FreeVC[58] 0.653 +0.31% 0.233 -16.19% 0.082 -20.39%
Table V: SER performance and fairness under different combinations of BS and VC. Subscripts indicate relative improvements over the ERM baseline. ✓ indicates the component is included.
# BS VC Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$ $\downarrow$ $\boldsymbol{\mathrm{DP}_{\text{gap}}}$ $\downarrow$
1 0.659 +1.23% 0.259 -6.83% 0.097 -5.83%
2 0.652 +0.15% 0.240 -13.67% 0.088 -14.56%
3 0.653 +0.31% 0.233 -16.19% 0.082 -20.39%
@@ VII-D Effect of Bias Selection and Voice Conversion
We compared three configurations that vary in the use of bias selection (BS) and voice conversion (VC) augmentation. Table V reports their relative improvement over the ERM baseline, which does not apply BS or VC.
  • Configuration#1 included BS but not VC. In this setting, bias-contrary samples were assigned higher weights during training, but no data were augmented.
  • Configuration#2 included VC but not BS. VC was applied to randomly selected pairs of utterances without considering their bias status.
  • Configuration#3, which represents our complete method, applied both BS and VC by selecting bias-guiding and bias-contrary samples and generating converted examples accordingly.
As shown in Table V, our method (#3) achieved the best overall improvement compared to ERM, striking the best balance between classification performance and fairness. It produced the largest reductions in $\mathrm{TPR}_{\text{gap}}$ and $\mathrm{DP}_{\text{gap}}$, along with competitive Macro F1. Configuration#1 improved the accuracy slightly but offered a limited fairness improvement. Configuration#2 yielded more fairness improvement than #1, but its overall results remained inferior to ours.